### v2

# Visual Odometry with Transformers

Abstract
Despite the rapid development of large 3D models, classical optimization-based approaches dominate the field of visual odometry (VO). Thus, current approaches to VO heavily rely on camera parameters and many hand-crafted components, most of which involve complex bundle adjustment and feature matching processes. Although disregarded in the literature, we find it problematic in terms of both (1) speed, that performs bundle adjustment requires a significant amount of time; and (2) scalability, as hand-crafted components struggle to learn from large-scale training data. In this work, we introduce a simple yet efficient architecture, Visual odometry Transformer (VoT), that formulates monocular visual odometry as a direct relative pose regression problem. Our approach streamlines the monocular visual odometry pipeline in an end-to-end manner, effectively eliminating the need for handcrafted components such as bundle adjustment, feature matching, or camera calibration. We show that VoT is up to $\sim 4$ times faster than traditional approaches, yet with competitive or better performance. Compared to recent 3D foundation models, VoT runs $\sim 10$ times faster with strong scaling behavior in terms of both model sizes and training data. Moreover, VoT generalizes well in both low-data regimes and previously unseen scenarios, reducing the gap between optimization-based and end-to-end approaches. Code will be released.
1 Introduction
[…]
The goal of monocular visual odometry is to estimate a camera’s position and orientation from a sequence of video frames [cadena2016past_present_future_of_slam]. In recent years, monocular visual odometry has gained increasing attention across various fields, including augmented reality, virtual reality, autonomous driving, and robotics. This growing interest is particularly notable when compared to systems that rely on stereo vision [wang2017stereo_dso, engel2014lsd] or multimodal inputs, such as visual-inertial odometry [stumberg2018direct_sparse_visual_inertial, christian2015robotics_science_systems]. Although monocular setups present greater challenges, their simplicity in deployment enables even broader applicability in real-world scenarios. Therefore, this paper focuses on monocular visual odometry using neural network-based approaches.
There has been rapid progress in monocular visual odometry over a short period, highlighting the potential of learning-based approaches. The dominant approach brought by this progress utilizes learnable features as well as differentiable optimization layers within deep learning frameworks [chen2024leapvo, teed2023dpvo]. While these designs improve convergence and robustness in real-world conditions, their performance critically relies on post-processing techniques, such as bundle adjustment or feature matching, to refine the estimates of camera poses. However, the limitations are still evident in real-world applications as the prediction in the wild has to undergo a slow post-processing process. Furthermore, hand-crafted elements incorporate prior domain knowledge, making it hard to scale with training data.
To this end, we shift our attention to the fast and scalable architecture. Recent successes in machine translation [vaswani2017transformer], image recognition [dosovitskiy2020vit], and object detection [carion2020detr] demonstrated that a simple transformer is effective enough for end-to-end structured prediction. By learning a mapping directly from video input to camera poses, the model becomes less dependent on predefined priors and can better capture complex temporal and spatial relationships. This design choice aligns with recent evidence [nguyen2025pit], suggesting that reducing inductive bias enables models to generalize more effectively from data. Most importantly, the end-to-end formulation enables direct prediction of camera trajectories without relying on heavy post-processing.
In this paper, we introduce the Visual odometry Transformer (VoT), a fully end-to-end framework for monocular visual odometry. Our approach formulates visual odometry as a direct pose prediction task. It mainly differs from previous models in the use of encoder-decoder architecture. The encoder is initialized with a pre-trained backbone to extract image features, while the decoder employs both temporal and spatial attention to model interactions across frames. The decouple between feature extraction with encoder and feature interaction with decoder significantly boosts the inference speed and scalability of the model. Fig. 1 shows that our end-to-end model is around ten times faster than 3D foundation models and at least three times than those with hand-crafted components while exhibiting similar or even better performance on multiple benchmarks.
Our key contributions can be summarized as follows:
  • VoT is the simple architecture by far for visual odometry as it commissions the pre-trained transformer as encoder and a light-weight transformer as decoder. This design inherently leads to significant speedup in our runtime while still being scalable in terms of both model sizes and training data;
  • For the first time, we achieve competent performance on several benchmarks of visual odometry without using any hand-crafted components and simply predicting camera poses in an end-to-end manner;
  • We showcase that end-to-end learning with VoT is beneficial especially when large-scale train data is available while it still generalizes well to low-regime data and unseen scenarios.
2 Related Work
Visual odometry. Visual odometry systems estimate the position and orientation of a camera using video input. Unlike SLAM, which corrects errors via loop closure [cadena2016past_present_future_of_slam, campos2021orbslam3, yugay2024magicslam], these systems have to cope with camera tracking errors (drift). Many different modalities of visual odometry have been explored by past work, including visual-inertial odometry [christian2015robotics_science_systems, stumberg2018direct_sparse_visual_inertial] and stereo visual odometry [engel2014lsd, wang2017stereo_dso]. While improving the overall performance, the multi-modal setting limits the potential to scale up the system with training data. Here, we focus on the monocular case, where the only input is a monocular video stream. Traditional monocular visual odometry methods [engel2014lsd, engel2018direct, campos2021orbslam3] are sensitive to illumination and rolling shutter artifacts, and more importantly, cannot adequately estimate the scale of the scene. To overcome these limitations, VoT leverages a large pre-trained encoder and extensive training data for robust generalization.
Figure 2: VoT architecture. Given multiple input frames, a frozen image encoder extracts per-image token embeddings. Camera embeddings are then concatenated to aggregate the information for camera pose estimation. The embeddings are decoded by $L$ repeating decoder blocks with temporal and spatial attention modules. The rotations are projected onto the SO(3) manifold to ensure valid relative rotations.
Deep monocular visual odometry. Deep learning has further advanced monocular visual odometry in both supervised [wang2017deepvo, tartanvo2020tartanVO, teed2020deepv2d, teed2023dpvo] and unsupervised [yin2018geonet, ranjan2019compeptitivecollaboration, alisha2019unsupervised_learning_of_depth_and_ego_motion, li2020selfsuperviseddeepvisualodometry] settings. DeepVO [wang2017deepvo] uses recurrent networks for temporal modeling, while SfMLearner [alisha2019unsupervised_learning_of_depth_and_ego_motion] jointly learns depth and motion without labels. DPVO [teed2023dpvo], inspired by RAFT [zachary2020raft], combines flow, confidence, and geometry with iterative GRU updates, while LeapVO [chen2024leapvo] increases robustness in dynamic scenes via keypoints. While being great at optimizing the rotations via bundle adjustment, these methods fail to estimate the scale of the scene, making their applicability in real-world scenarios limited. Moreover, they require camera parameters and are slow, which makes their deployment difficult. In contrast, VoT directly regresses relative camera poses in metric space and does not require camera parameters.
Prior work introduced TSFormer [francani2025tsformer], which predicts camera poses directly from video sequences using an end-to-end approach. However, its performance remains significantly below that of hand-crafted methods [teed2023dpvo, chen2024leapvo] due to architecture, small training dataset, rotation representation, and inference scheme. Existing end-to-end methods struggle to perform in real-world settings because they rely on small, hard-to-scale architectures, and the gap between them and traditional methods remains substantial. By comparison, VoT uses a scalable end-to-end design which provided enough data to outperform traditional methods. When the data is limited, our method significantly reduces the performance gap between end-to-end and optimization-based pipelines, while being significantly faster.
Video transformers. Since their introduction in machine translation [vaswani2017transformer], transformers have become the dominant architecture in NLP [devlin2019bert, openai2023gpt4] and vision [he2022mae, oquab2024dinov2]. Central to this framework, self-attention aggregates information across entire sequences, but its quadratic complexity makes direct application to videos inefficient. To mitigate this, prior work has proposed factorizing 3D attention into spatial and temporal components [zhang2021vidtr] or employing efficient variants such as divided space-time, sparse local-global, and axial attention [bertasius2021timesformer, arnab2021vivit, zhao2022tuber]. In this work, we adopt the principle of efficient video learning by decoupling self-attention into separate temporal and spatial attention components.
Large 3D models. A trend in 3D vision is the development of large-scale models that jointly estimate camera poses and dense geometry [wang20243d_reconstruction_with_spatial_memory, wang2025vggt, wang2025cut3r], or learn versatile representations for downstream tasks [wang2024dust3r, leroy2024mast3r]. These models are increasingly viewed as foundational, yet despite strong task-agnostic results they suffer from drift on long video sequences and produce scale-ambiguous poses. In contrast, VoT focuses solely on the task of camera pose estimation, showing improved accuracy and less drift. Moreover, our model is significantly faster due to its size.
3 Method
In this section, we describe the architecture of VoT; see Fig. 2 for its overview. It consists of three main components: a pre-trained encoder to extract a compact representation of each frame, a transformer-based decoder, and a small feed-forward network that makes the final pose predictions. Unlike previous methods such as DVPO [teed2023dpvo] or LeapVO [chen2024leapvo], VoT can be implemented in any deep learning framework, thanks to its streamlined architecture. Following the spirit of end-to-end frameworks [carion2020detr], our inference code can be written purely in PyTorch, without additional complexities. We believe that the simplicity and accessibility of our approach will encourage further participation from the research community.
3.1 Architecture
Pre-trained encoder. We utilize pre-trained networks commonly employed in previous studies [wang2024dust3r, wang2025cut3r] for encoding and extracting features from video frames. As the Vision Transformer (ViT) [dosovitskiy2020vit] has become the dominant architecture in computer vision, our pre-trained encoders are based on ViT. Specifically, we denote the sequence of input video frames as $V\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ is the number of frames in the video and $H,W$ denote the resolution of each frame. Each input image is then tokenized into $(h\cdot w)$ non-overlapping patches with $h{=}\frac{H}{p},w{=}\frac{W}{p}$ and $p$ as patch size. These tokens are processed through a series of ViT layers, producing image features $F\in\mathbb{R}^{T\times(h\cdot w)\times d}$, where $d$ is the hidden dimension of the ViT. The ViT layer follows the standard architecture, consisting of a multi-head self-attention module and a feed-forward network. Since the transformer architecture is permutation-invariant, we augment it with sinusoidal position encodings [vaswani2017transformer].
Time-space decoder. The decoder consists of a stack of $L$ identical layers, each containing three sub-layers. The first sub-layer is a multi-head temporal attention module, followed by a multi-head spatial attention module, and the final sub-layer is a feed-forward network. Unlike standard transformer layers that only use self-attention, the interleaving of temporal and spatial attention within the decoder layers enables each feature in a single frame to efficiently capture long-range dependencies with other features. This ability to learn long-context information effectively is crucial when processing a large number of frames. As we demonstrate in our experiments, using a larger number of frames is key to achieving accurate camera pose predictions. To summarize the information in each frame, we use camera embeddings, which are learnable embeddings in spatial attention sub-layers.
We start by concatenating a camera embedding to each of the image features, resulting in input features to the decoder $F_{0}{=}[\mathrm{ce},F]\in\mathbb{R}^{T\times(h\cdot w+1)\times d}$ where $\mathrm{ce}\in\mathbb{R}^{T\times 1\times d}$ indicate the camera embeddings. We denote the inputs to the $(n+1)^{\text{th}}$ decoder layer by $F_{n}\in\mathbb{R}^{T\times(h\cdot w+1)\times d}$. The $(n+1)^{\text{th}}$ decoder layer then outputs $F_{n+1}\in\mathbb{R}^{T\times(h\cdot w+1)\times d}$ of the same size. Specifically, the temporal attention performs the scaled dot-product attention in the $i$-th head along the temporal dimension as [^1] The notation [:, 1:, : ] indicates the submatrix that removes the first column, as in Numpy.:
$\begin{split}\hat{Q}=\hat{K}=\hat{V}&=F\text{[:, 1:, :]}^{\top},\\ \hat{Q},\hat{K},\hat{V}&\in\mathbb{R}^{(h\cdot w)\times T\times d}\end{split}$ (1)
$\begin{split}\hat{\mathrm{head}}_{i}&=\mathrm{Attention}(\hat{Q}W^{\hat{Q}}_{i},\hat{K}W^{\hat{K}}_{i},\hat{V}W^{\hat{V}}_{i}),\\ \hat{\mathrm{head}}_{i}&\in\mathbb{R}^{(h\cdot w)\times T\times d_{h}},\end{split}$ (2)
where $W^{\hat{Q}}_{i},W^{\hat{K}}_{i},W^{\hat{V}}_{i}\in\mathbb{R}^{d\times d_{h}}$ are the learned projection matrices for query, key, and value. Here, we omit the layer index by treating $F_{n}=F$.
[…]
$\begin{split}\hat{F}&=\mathrm{Concat}(\hat{\mathrm{head}}_{i},\ldots,\hat{\mathrm{head}}_{h})^{\top}W^{\hat{O}},\\ \hat{F}&\in\mathbb{R}^{T\times(h\cdot w)\times d}\end{split}$ (3)
$\begin{split}\mathrm{TemporalAttention}(\hat{Q},\hat{K},\hat{V})&=[F\text{[:,:1,:]},\hat{F}],\\ \mathrm{TemporalAttention}&\in\mathbb{R}^{T\times(h\cdot w+1)\times d}\end{split}$ (4)
where $W^{\hat{O}}_{i}\in\mathbb{R}^{d_{h}\times d}$ is the output projection and $F\text{[:,:1,:]}\in\mathbb{R}^{T\times 1\times d}$ indicates the camera embeddings copied from $F$.
Following the temporal attention, we apply the spatial attention along the spatial dimension, including the camera embedding as:
$\bar{Q}=\bar{K}=\bar{V}=\hat{F}\in\mathbb{R}^{T\times(h\cdot w+1)\times d_{h}}\\$ (5)
$\begin{split}\bar{\mathrm{head}}_{i}&=\mathrm{Attention}(\bar{Q}W^{\bar{Q}}_{i},\bar{K}W^{\bar{K}}_{i},\bar{V}W^{\bar{V}}_{i})\\ &\in\mathbb{R}^{T\times(h\cdot w+1)\times d_{h}},\end{split}$ (6)
where $W^{\bar{Q}}_{i},W^{\bar{K}}_{i},W^{\bar{V}}_{i}\in\mathbb{R}^{d\times d_{h}}$ are the learned projection matrices for query, key, and value.
Similarly, the multi-head spatial attention aggregates $\bar{\mathrm{head}}_{\{1...h\}}$ together.
$\begin{split}\mathrm{SpatialAttention}(\bar{Q},\bar{K},\bar{V})&=\mathrm{Concat}(\bar{\mathrm{head}}_{i},\ldots,\bar{\mathrm{head}}_{h})W^{\bar{O}}\\ &\in\mathbb{R}^{T\times(h\cdot w+1)\times d}\end{split}$ (7)
where $W^{\bar{O}}_{i}\in\mathbb{R}^{d_{h}\times d}$ is the output projection.
Ultimately, the camera embedding, $e\in\mathbb{R}^{T\times d}$, will be utilized to predict the camera poses for each frame.
Figure 3: Attention maps from the VoT decoder. Each row shows an original image with a selected query (red square), followed by attention maps from the four subsequent frames. To estimate relative camera pose, VoT attends to the related image regions, resembling the behavior of classical keypoint-based odometry methods.
3.2 Camera Poses Prediction
To regress the relative camera poses, the regression head takes a camera embedding from each frame as its input. A single linear layer then projects the embeddings to the rotation $\mathbb{F}_{R}\in\mathbb{R}^{(T-1)\times 9}$ and translation $\mathbb{F}_{t}\in\mathbb{R}^{(T-1)\times 3}$ vectors. Rotation vectors are projected to the closest valid rotation matrix by solving the special orthogonal Procrustes [bregier2021deepregression] problem:
$\text{Procrustes}(\mathbb{F}_{R})={\arg\min}_{\hat{\mathbb{R}}\in SO(3)}\|\hat{\mathbb{R}}-\mathbb{F}_{R}\|_{F}^{2}.$
The optimization problem is solved via singular value decomposition of $\mathbb{F}_{R}$ following [umeyama1991lse]. $\mathbb{F}_{t}\in\mathbb{R}^{(T-1)\times 3}$ is used as a translation prediction directly.
[…]
$\mathcal{L}_{\text{rotation}}(R,\hat{R})=\cos^{-1}\left(\frac{\text{Tr}(R^{\top}\hat{R})-1}{2}\right),$ (8)
and the translation loss, denoted as $\mathcal{L}_{\text{translation}}$, is the L1 loss between the predicted translation $\hat{t}$ and the ground-truth translation $t$:
$\mathcal{L}_{\text{translation}}(t,\hat{t})=\|t-\hat{t}\|_{1}.$ (9)
The final loss for VoT training is a weighted loss between the rotation and translation losses:
$\mathcal{L}=\lambda\cdot\mathcal{L}_{\text{rotation}}+\gamma\cdot\mathcal{L}_{\text{translation}}.$ (10)
4 Experiments
Datasets. Our primary training leverages ARKitScenes [baruch2022arkit], a large-scale dataset, providing rich and diverse indoor scenes. We further incorporate ScanNet [dai2017scannet] to broaden indoor coverage and KITTI [geiger2012kitti] to expose the model to outdoor pose distributions. Out-of-distribution performance is evaluated on TUM_RGBD [sturm2012tumrgbd] which is not included in the training set. ARKitScenes [baruch2022arkit] is a large-scale indoor dataset collected with LiDAR-equipped mobile devices, comprising over 5,000 sequences from 1,661 environments and totaling more than 11 million frames, making it one of the most comprehensive resources for indoor 3D scene understanding. To complement it, ScanNet [dai2017scannet] provides over 2 million frames across 1,613 scenes. For outdoor settings, KITTI [geiger2012kitti] offers RGB images and ground-truth poses collected with GPS and LIDAR. Finally, TUM_RGBD [sturm2012tumrgbd] features indoor recordings with motion-capture ground truth. The datasets exhibit variation in camera calibration, poses, image resolution, as well as real-world artifacts such as rolling shutter, motion blur, and lighting variation.
Metrics. We evaluate visual odometry using Absolute Translation Error (ATE) and Absolute Rotation Error (ARE) [teed2023dpvo, chen2024leapvo]. ATE is the RMSE of translation between estimated and ground-truth trajectories, while ARE is the RMSE of angular orientation differences. In addition, we report Relative Translation Error (RTE) and Relative Rotation Error (RRE), which measure the drift in translation and rotation over unit trajectory segments. In the tables, results are coded best and $\underline{\text{second best}}$.
A note on alignment. While it is common practice to report evaluation metrics after applying rigid alignment and scale correction to the predicted trajectory, we believe this practice can be impractical for real-world applications, where ground-truth is unavailable. To take this real-world constrain into account we report both aligned and unaligned metrics.
Figure 4: Visualized Predicted Trajectories from Test Splits of ScanNet and ARKit Datasets. Most methods cannot predict a coherent trajectory without aligning to the ground truth. In contrast, VoT predicts an accurate trajectory without alignment.
Baselines. We compare our method against state-of-the-art visual odometry models [teed2023dpvo, chen2024leapvo]. For DPVO [teed2023dpvo], results are averaged over three runs with different random seeds to account for variability. We further benchmark against recent large-scale 3D models that directly predict camera poses [wang2025vggt, wang20243d_reconstruction_with_spatial_memory]. Because these models cannot process long video sequences in a single pass, inputs are split into temporally continuous chunks, with maximum lengths of 30 and 90 frames on our GPUs, respectively; the predicted poses are then concatenated to form the full trajectory. We also compare with a classical [campos2021orbslam3] and a state-of-the-art [murai2024mast3rslam] monocular SLAM systems, both using only RGB input. For fairness, loop closure is disabled during evaluation. Finally, we include a recent end-to-end visual odometry model [francani2025tsformer] in our comparisons.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.08 133.92 0.03 4.01
DPVO [teed2023dpvo]†∗ 0.49 20.01 0.02 0.49
LeapVO [chen2024leapvo]†∗ 0.89 58.42 0.04 1.10
TSFormer [francani2025tsformer] 141.34 56.32 0.37 0.62
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.67 82.61 0.09 4.86
VGGT [wang2025vggt] 2.26 124.64 0.16 4.81
Mast3r-SLAM-VO [murai2024mast3rslam] 0.28 25.49 0.19 17.89
VoT (Ours) 0.25 10.89 0.01 0.44
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 2.58 111.58 0.17 4.01
DPVO [teed2023dpvo]†∗ 5.48 18.68 0.04 0.48
LeapVO [chen2024leapvo]†∗ 28.31 49.25 0.62 1.09
TSFormer [francani2025tsformer] 285.22 97.53 0.52 0.62
CUT3R [wang20243d_reconstruction_with_spatial_memory] 2.42 97.70 0.12 4.86
VGGT [wang2025vggt] 2.94 92.37 0.15 4.81
Mast3r-SLAM-VO [murai2024mast3rslam] 0.60 14.40 0.36 17.90
VoT (Ours) 0.51 13.40 0.01 0.44
Table 1: Pose estimation metrics on ARKitScenes indoor settings. Methods marked with use bundle adjustment and methods marked with use camera parameters. Provided enough training data, VoT shows superior performance.
Implementation details. We use a frozen CroCo [philippe2022croco] backbone trained within the DUST3R [wang2024dust3r] framework, consisting of 300 million parameters. We employ 12 alternating time-space attention blocks, totalling 200 million parameters, for the decoder. Our model is trained with the AdamW [loshchilov2017adamw] optimizer for 300 epochs. We adopt a cosine learning rate schedule with an initial learning rate of 0.00001 and a warmup phase of 30 epochs. Our model takes 8 input views, sampled at intervals of 3 frames. Input frames are resized to a resolution of $\times224$. Training runs on 12 NVIDIA RTX H100 GPUs for 5 days.
4.1 Visual Odometry Results
VoT is competitive without bells and whistles. As shown in Tabs. 1, 2 and 3, VoT achieves competitive performance on both indoor and outdoor datasets, despite not relying on hand-crafted components. This is particularly notable in real-world settings, where unaligned trajectory metrics are more realistic due to the lack of ground-truth poses.
Large 3D models such as CUT3R [wang20243d_reconstruction_with_spatial_memory] and VGGT [wang2025vggt], although trained on diverse multi-task 3D datasets - including both ScanNet and ARKitScenes - fail to generalize effectively. CUT3R and VGGT exhibit substantial drift in long sequences, while Mast3r-SLAM-VO [murai2024mast3rslam] suffers from scale ambiguity and sparse predictions, resulting in degraded ATE and failure on the KITTI dataset.
Among odometry-specific methods, VoT exhibits strong generalization across datasets. In contrast, LeapVO [chen2024leapvo], trained solely on synthetic data, struggles on real-world videos. DPVO [teed2023dpvo] achieves competitive results on ScanNet and KITTI but fails to accurately estimate the scale (ATE). ORB-SLAM3 [campos2021orbslam3] fails to reliably estimate scale, leading to poor performance even with loop closure and bundle adjustment enabled. TSFormer [francani2025tsformer], trained on a small dataset, without strong encoders or robust rotation representations, fails to predict accurate trajectories.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.07 141.78 0.04 3.75
DPVO [teed2023dpvo]†∗ 0.16 8.58 0.02 0.34
LeapVO [chen2024leapvo]†∗ 0.74 72.78 0.02 0.43
TSFormer [francani2025tsformer] 105.21 61.42 0.32 0.45
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.38 24.13 0.02 0.65
VGGT [wang2025vggt] 1.08 80.84 0.07 1.42
Mast3r-SLAM-VO [murai2024mast3rslam] 0.22 19.31 0.22 6.59
VoT (Ours) 0.15 8.50 0.01 0.42
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.91 100.42 0.04 3.75
DPVO [teed2023dpvo]†∗ 1.75 5.91 0.02 0.34
LeapVO [chen2024leapvo]†∗ 10.84 43.60 0.09 0.43
TSFormer [francani2025tsformer] 285.22 97.53 0.52 0.62
CUT3R [wang20243d_reconstruction_with_spatial_memory] 4.85 135.52 0.02 0.65
VGGT [wang2025vggt] 1.56 31.73 0.06 1.42
Mast3r-SLAM-VO [murai2024mast3rslam] 0.99 8.40 0.58 6.96
VoT (Ours) 0.34 8.35 0.01 0.42
Table 2: Pose estimation metrics on ScanNet indoor settings. Methods marked with use bundle adjustment and methods marked with use camera parameters. Without bells and whistles, VoT shows competitive performance.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 80.81 157.31 2.82 1.29
DPVO [teed2023dpvo]†∗ 9.74 13.63 0.93 0.06
LeapVO [chen2024leapvo]†∗ 48.86 67.98 0.39 0.96
TSFormer [francani2025tsformer] 15.77 93.57 0.45 0.74
CUT3R [wang20243d_reconstruction_with_spatial_memory] 17.92 73.76 0.50 0.58
VGGT [wang2025vggt] 45.07 98.23 1.78 0.51
Mast3r-SLAM-VO [murai2024mast3rslam]
VoT (Ours) 9.66 31.12 0.32 0.42
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 217.06 51.23 1.05 1.29
DPVO [teed2023dpvo]†∗ 194.55 0.94 0.85 0.06
LeapVO [chen2024leapvo]†∗ 211.80 39.78 0.94 0.96
TSFormer [francani2025tsformer] 140.10 28.45 0.78 0.74
CUT3R [wang20243d_reconstruction_with_spatial_memory] 112.36 22.07 0.71 0.58
VGGT [wang2025vggt] 205.67 16.32 2.13 0.51
Mast3r-SLAM-VO [murai2024mast3rslam]
VoT (Ours) 63.10 11.69 0.65 0.42
Table 3: Pose estimation metrics on KITTI outdoor settings. Methods marked with use bundle adjustment and methods marked with use camera parameters. ✗ denotes the methods that failed on some of the sequences. Even though KITTI training data amounts to only several videos, VoT shows competitive performance.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 0.80 147.39 0.05 3.16
DPVO [teed2023dpvo]†∗ 0.10 3.08 0.04 0.44
LeapVO [chen2024leapvo]†∗
TSFormer [francani2025tsformer] 0.82 153.35 0.05 0.92
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.87 29.42 0.02 1.41
VGGT [wang2025vggt] 1.05 70.34 0.03 0.77
Mast3r-SLAM-VO [murai2024mast3rslam] 0.83 150.51 1.36 60.03
VoT (Ours) 0.21 7.91 0.01 0.65
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.65 76.14 0.03 3.16
DPVO [teed2023dpvo]†∗ 0.94 3.01 0.07 0.44
LeapVO [chen2024leapvo]†∗
TSFormer [francani2025tsformer] 200.87 105.86 0.58 0.92
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.51 60.86 0.02 1.41
VGGT [wang2025vggt] 1.61 32.66 0.03 0.77
Mast3r-SLAM-VO [murai2024mast3rslam] 1.21 47.81 1.32 60.03
VoT (Ours) 0.47 10.37 0.01 0.65
Table 4: Pose estimation metrics on TUM_RGBD out-of-distribution setting. Methods marked with use bundle adjustment and methods marked with use camera parameters. ✗denotes the methods that failed on some of the sequences. VoT performs well on unaligned metrics but lags behind in aligned settings.
4.2 Ablation Studies
We conduct ablation studies to assess key design choices and analyze scaling trends with respect to model capacity and dataset size. For architectural ablations (e.g., attention mechanisms, backbone, and rotation representation), we use two input views with a stride of 10 and train for 150 epochs. For scaling experiments (e.g., varying dataset size), models are trained to convergence for 300 epochs to reflect scalability accurately. All ablations are evaluated on the ScanNet [dai2017scannet] dataset.
VoT delivers real-time prediction. To compare the speed between approaches, we measure their runtime on ScanNet [dai2017scannet] scene0000_00 using the same environment. As shown in Fig. 1, VoT achieves a considerably faster runtime, more than $3\times$ speedup, compared to existing methods. This efficiency stems from its compact, end-to-end architecture, in contrast to larger 3D models. Moreover, unlike LeapVO, DPVO, and Mast3R-SLAM-VO, VoT does not rely on additional processing steps or bundle adjustment, contributing to both its speed and strong performance.
VoT is competitive on unseen datasets and camera parameters. Tab. 4 illustrates how VoT performs on a dataset not seen during training. Moreover, the camera parameters of the dataset are drastically different from those in the training set. While DPVO leverages effective bundle-adjustment for superior performance, VoT demonstrates competitive results and robust generalization compared to other methods.
(a) Scaling behavior w.r.t. training data
(b) Scaling behavior w.r.t. model capacity
Figure 5: Scaling behavior of VoT. As the model scales in (a) training data (proportion of ARKitScenes data added to ScanNet) and (b) model capacity (number of decoder layers), absolute translation and rotation errors decrease. This suggests that VoT exhibits robust scaling behavior.
VoT exhibits robust scaling behavior. Fig. 5 illustrates how VoT scales with respect to training data size and model capacity. VoT shows consistent performance improvements with both increased training data and a larger number of trainable parameters. Due to computational limits, we cap the number of decoder layers at 12. We analyze the scaling behavior with respect to encoder size and number of input views in the Supplementary Material.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
MAE 1.26 28.30 0.08 2.22
DinoV2 1.05 31.52 0.07 2.09
DinoV2* 0.96 29.06 0.07 1.95
CroCoV2 1.02 32.58 0.07 2.46
CroCoV2* 0.55 15.09 0.04 1.25
Table 5: Backbone ablation. CroCoV2* denotes the model trained within the Dust3r [wang2024dust3r] framework, and DinoV2* denotes the backbone train in VGGT framework [wang2025vggt]. The results highlight the critical role of the architecture of the backbone, with geometry-focused supervision yielding better results across all metrics.
Attention Type ATE[m] $\downarrow$ ARE[deg] $\downarrow$ RTE[m] $\downarrow$ RRE[deg] $\downarrow$ GFLOPs $\downarrow$
Full Attention 1.57 60.24 0.09 4.04 380
Time-Space Attention 1.27 35.68 0.08 2.22 163
Table 6: Ablation of attention mechanisms. Time-space attention outperforms full attention while being considerably more efficient. We attribute this to its ability to capture temporal changes by linking spatially corresponding patches across frames.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Euler Angles 0.78 20.64 0.05 1.47
Quaternion 0.68 18.92 0.05 1.44
Plucker Rays [zhang2024raydiffusion] 0.66 16.11 0.05 1.48
Rotation Matrix 0.55 15.09 0.04 1.25
Table 7: Ablation of rotation representations. Projecting predictions onto the nearest valid rotation matrix under the Frobenius norm yields the best performance.
Architecture is the key determinant of backbone performance. We compare features extracted from various pre-trained backbones in Tab. 5. While DINOv2 [oquab2024dinov2] and CroCoV2 [philippe2022croco] yield comparable results, the CroCoV2 encoder trained within the Dust3R [wang2024dust3r] framework that uses a geometry-focused architectur- achieves substantially better performance (i.e., approximately $2\times$ lower error across all metrics). Moreover, CroCoV2 trained within Dust3R framework shows superior performance to DinoV2 backbone trained on comparable amounts of data [wang2025vggt] and similar tasks. This suggests that the backbone architecture is the key components for a strong performance for our method.
SO(3) projection enhances rotation accuracy. In Tab. 7, we evaluate different rotation representations. Projecting outputs onto the nearest valid rotation matrix on the SO(3) manifold using the Frobenius norm consistently yields the best results, highlighting its suitability for our formulation.
Time-space attention achieves superior performance with lower computational cost. Tab. 6 compares full self-attention and time-space attention. Time-space attention not only reduces computational cost but also consistently improves accuracy. We attribute this to its matching-like spatial mechanism (see Fig. 3), where patches at corresponding locations across frames exchange temporal information, enhancing motion-aware feature encoding.
Limitations and future work. We do not claim VoT can generalize to all the datasets. Since it is trained on mostly static environments, performance may be limited in dynamic settings. Future improvements could come from scaling to more diverse datasets collected with different devices and careful calibration, as well as expanding coverage to a wider variety of scenes. Employing systematic data curation and incorporating larger models or advanced pre-trained components are also promising directions that could enhance both generalization and overall performance.
5 Conclusion
We presented VoT, an end-to-end method for visual odometry systems based on transformers for direct relative pose prediction. The approach achieves strong results compared to modern state-of-the-art odometry pipelines and large 3D models on the challenging indoor and outdoor datasets. VoT has a flexible architecture easily extensible to various backbones, with competitive results. Importantly, it can generalize to out-of-distribution camera parameters and image data. It exhibits strong scaling behavior, indicating that its performance could be further enhanced with sufficient data and computational resources. Finally, it is considerably faster than alternatives, being more compact than large 3D models, and, unlike other odometry systems, it does not require test-time optimization. We anticipate that future work will further advance end-to-end odometry systems.
Acknowledgements
This work has been financially supported by TomTom, the University of Amsterdam, and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
[…]
Supplementary Material
6 Appendix
We test how well VoT scales with different backbone sizes and numbers of input views.
Performance of VoT with number of input views. Figure 6 illustrates how VoT, trained on the full dataset, scales with the number of input views. While the performance plateaus in terms of absolute metrics, it consistently improves in relative metrics. Due to computational constraints, the number of input views in our experiments is limited to a maximum of 12.
Figure 6: Performance of VoT with number of input views on ScanNet [dai2017scannet] dataset. As the number of input views increases, absolute translation and rotation errors initially decrease before plateauing. In contrast, relative metrics continue to improve steadily, highlighting the potential of VoT when using a long-range input sequence of views.
Performance of VoT with backbone size. In Fig. 7, we show how the performance of VoT varies with the size of the backbone. Since CroCO [philippe2022croco] is only available in a single size, we perform ablation using Dino-V2 [oquab2024dinov2] backbones. We observe that performance improves as the size of the backbone is increased. Due to computational limits, we cap the backbone size at DinoV2 [oquab2024dinov2] large.
Figure 7: Performance of VoT with backbone size on ScanNet [dai2017scannet] dataset. As the number of backbone parameters increases, the model’s performance improves.